OpenStudy (anonymous):

How many solutions does the system of equations have? 2x + 3y = 6 y= -2/3x-1 A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

First, let's put them both in the same form, in this case, we'll do slope-intercept form. 2x-2x+3y=6-2x 3y=-2x+6 3y/3=-2x/3+6/3 y=-2/3x+2 So our two equations are: y=-2/3x+2 and y=-2/3x-1 The identical slopes tell us that these two lines are parallel, which means they will never intersect. Therefore, your answer is 0.